Subject: 3.5-m status report


Submitted: Mon, 11 Sep 1995 09:43:39 -0600

Message number: 15 (previous: 14, next: 16 up: Index)

             APO 3.5-m Telescope, Instruments, and related systems

                           Status: 10 September 1995


Attention was paid to the various telescope and enclosure drives this past
month.  A problem with the azimuth drives was solved by changing a
parameter in the controller, eliminating frequent motor stoppages due to
overcurrent. The altitude drive continues to have an intermittent
offsetting problem, which is under investigation. The more recent pointing
models show ~3 arcsec rms blind pointing, but actual on-sky performance is
sometimes inconsistent.  The enclosure drives continue to lose the tuning
of their servo adjustments on a weekly basis--an engineer from the company
that built the enclosure drives is being brought to the site to make a
diagnosis.  Nasmyth rotator stalling problem was solved by installing a new
stronger motor.  The primary and secondary mirrors were both recollimated
during engineering time. Image quality continues to be tested, and recent
data show slight but noticeable astigmatism in the Nasmyth images which has
not been seen before--this is to be confirmed through a Hartmann test this
coming week.  Also, since instrumental/system throughput has been under
question, the reflectivity of the three mirrors were checked, showing a net
reflectivity of the telescope of ~60%, nominal being ~73% for fresh
coatings.  Many people too numerous to list helped with the testing and
repairs of telescope, including site staff but also many research staff and
students from the ARC member institutions.

As a case in point, Ed Kibblewhite and his students from UChicago analyzed
recent data taken with their adaptive optics package (ChAOS), and
discovered a strong image motion resonance at 20Hz.  Not knowing if this
signal was internal to their instrument or in the image itself, tests of
the telescope were performed which measured image motion with the ChAOS
package, plus monitoring mechanical motions of the telescope itself.  The
20 Hz motions were found to be in the telescope, and are being driven by
the altitude velocity servo which updates its input to the motors at 20 Hz.
How this 20 Hz image motion effects the point spread function is presently
being tested, and we have some optimism that a straightforward fix may
correct this problem.

As reported last month, we found that the circumferential cracks in the
enclosure wheels had grown in the past year, and a campaign to replace
these wheels at the earliest opportunity has been launched.  One
replacement wheel has been ordered (made with a slightly different steel),
and in parallel we are exploring a new wheel design that has been suggested
by L&F which may be better, cheaper, and faster (or at least better and
cheaper).  Replacement of the worst cracked wheel is planned for late
October, but if any wheel crack accelerates or chunks break out of the
wheel before the planned repair, the telescope will have to be closed.

The experiment to test laser-cleaning of mirrors continues.

Routine measurements of r-naught and cloud scanner images of the sky
continue.  A recent all-sky "movie" of a dark night is available from our
web server.


Jon Brinkmann will take the post of Scientific Instruments Engineer,
arriving at APO this week after a brief swing through Chicago to talk with
the GRIM2 developers Mark Hereld and Bernie Rauscher.  Jon will initially
be tasked with projects related to instrument calibration systems, the
guider, and telescope baffling.

As promised, instrument reports have started to appear in the APO e-mail
exploder--see the APO web server for information on how to subscribe to the


Karen Gloria has written a detailed status report on DIS and placed it on
the e-mail exploder last week.  Of note, the bias/gain shift of the red
camera seems to have stopped since early summer, possibly as a result of
Richard Lucinio's attentions in June.  The question of DIS throughput is
being studied by several investigators.  Jim Gunn was at APO this past
weekend and attempted a DIS spectrophotometric measurement but was hampered
by clouds.  He has promised to send us a prescription for making this
measurement, which we will attempt during engineering time.  If other
scientists have proposals to measure DIS (or any instrument) performance
for the benefit of all, we will gladly consider those programs as
candidates for engineering time.


The instrument has been behaving very well.  Nothing to add to last month's


Pretty much unchanged since June, with the exception of some additional
tests and baffling which improves stray-light rejection of camera when
mounted at Nasmyth.


Return date to APO is TBD.  Not to be scheduled in upcoming quarter.


Return date to APO is TBD.  Not to be scheduled in upcoming quarter.


Closed-loop tests done in past week, results will be published.  New s/w
from vendor received which was supposed to fix run-time routines, compiled
but didn't execute.  Trying to get 512^2 chip at the site which has better
noise and sensitivity, and send 1024^2 camera back to vendor for overhaul.
Also trying to get Eric Deutsch's utility on-line for finding bright guide
stars for any target.


Used successfully for engineering studies of telescope vibrations and
measurement of wavefront.  Results to be available shortly.


Schedule for this project is TBD, but the initial capabilities were
expected to be ready by now, but have been delayed by repair work on
telescope and enclosure drives.

S/W and network-

Have had a few days of unreliable network access, again appears to be
related to routers in the Washington DC area.  A backup modem capability
has been implemented at the site, but to be fully useful requires certain s/w
and phone lines need to be installed at the remote sites.  Jim Fowler has
sent instructions to users, but no one appears to have yet implemented
modem capability at the remote observing rooms.

                                Bruce Gillespie

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 15 in the apo35-general archive. You can find
APO  the archive in /u/strauss/apo/mailer/apo35-general on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to